find sin(17pie/12)
Answer by lwsshak3(11628)

find exact value of sin(17pie/12)
***
sin(17π/12)
=sin(8π/12+9π/12)
=sin(2π/3+3π/4)
Identity: sin(x+y)=sinxcost+cosxsiny
sin(2π/3+3π/4)=sin(2π/3)cos(3π/4)+cos(2π/3)sin(3π/4)
=√3/2*-√2/2+-1/2*√2/2=-√6/4-√2/4=-(√6+√2)/4
check:
sin(17π/12)≈-0.9659 (with calculator)
Exact value as calculated=-(√6+√2)/4≈0.9659